I need help with another problem please help.
Martin leaves home at 9 a.m. bicycling at a rate of 24mi/hr.Two hours later,John leaves,driving at the rate of 48mi/hr.At what time will John cath up with Martin?

Answer by ankor@dixie-net.com(22740)

Martin leaves home at 9 a.m. bicycling at a rate of 24mi/hr. Two hours later,John leaves,driving at the rate of 48mi/hr.At what time will John catch up with Martin?
:
Let x = time (in hrs) that M is on the road
Then
(x-2) = time J is on the road, when he catches up
:
We know when this happens they will have traveled the same distance.
We can make a simple equationn from this fact: Dist = speed * time
:
J's dist = M's dist
48(x-2) = 24x
48x - 96 = 24x
48x - 24x = +96
24x = 96
x = 96/24
x = 4 hrs; from 9 am would be 1 pm when J catches up with M
:
:
Check solution using distance, J's time is 2hrs:
48(2) = 24(4)
:
Not that hard, right?